# Current affiliation:] Laboratoire Charles Coulomb (L2C), Univ. Montpellier, CNRS, Montpellier, France

### v1

@@ Abstract
We show that weakly nonreciprocal alignment leads to large-scale structure formation in flocking mixtures. By combining numerical simulations of a binary Vicsek model and the analysis of coarse-grained continuum equations, we demonstrate that nonreciprocity destabilizes the ordered phase formed by mutually aligning or anti-aligning species in a large part of the phase diagram. For aligning populations, this instability results in one species condensing in a single band that travels within a homogeneous liquid of the other species. When interactions are anti-aligning, both species self-assemble into polar clusters with large-scale chaotic dynamics. In both cases, the emergence of structures is accompanied by the demixing of the two species, despite the absence of repulsive interactions. Our theoretical analysis allows us to elucidate the origin of the instability, and show that it is generic to nonreciprocal flocks.
@@
When $\chi^{\textsl{{s}}\textsl{{u}}}=1$ for all $\textsl{{s}},\textsl{{u}}\in\{\textsl{{a}},\textsl{{b}}\}$, the model reduces to the classical VM, while other limiting cases such as fully anti-aligning mixtures [38], reciprocally anti-aligning populations ($\chi^{\textsl{{a}}\textsl{{a}},\textsl{{b}}\textsl{{b}}}>0$, $\bar{\chi}<0$, $\Delta\chi=0$[39, 40, 41] and self-anti-aligning species ($\chi^{\textsl{{a}}\textsl{{a}},\textsl{{b}}\textsl{{b}}}<0$) with aligning interspecies interactions ($\bar{\chi}>0$[42, 43] have been addressed in the literature. Here, we let particles of the same species align ($\chi^{\textsl{{a}}\textsl{{a}},\textsl{{b}}\textsl{{b}}}=1$) while varying the interspecies coulings $\chi^{\textsl{{a}}\textsl{{b}}}$ and $\chi^{\textsl{{b}}\textsl{{a}}}$. In the regime of strong nonreciprocity, where $|\Delta\chi|>|\bar{\chi}|$, we observe oscillatory behavior consistently with results reported in the literature [17, 31, 34]. Hereafter, however, we focus on the weakly nonreciprocal regime, where $|\Delta\chi|\leq|\bar{\chi}|$.
As we focus on the effects induced by nonreciprocity alone, we discard the possible influence of population heterogeneity [44] by setting $\eta^{\textsl{{a}},\textsl{{b}}}=\eta$ and $v_{0}^{\textsl{{a}},\textsl{{b}}}=v_{0}$  [^1] We show in [48] that our conclusions remain unaffected when introducing population heterogeneity.. We also consider equal densities $\rho^{\textsl{{a}},\textsl{{b}}}=\rho_{0}=N/L^{2}$, where $N$ denotes the size of each population and $L$ is the linear size of the square domain. Unless specified otherwise, we use $\rho_{0}=\tfrac{1}{2}$, $v_{0}=1$, $\bar{\chi}=\pm\tfrac{1}{2}$, and vary the angular noise $\eta$ and the nonreciprocity parameter $\Delta\chi$. Noting that the problem is symmetric under the exchange $\textsl{{a}}\leftrightarrow\textsl{{b}}$, we further restrict the analysis to $\Delta\chi\geq 0$. To construct the phase diagrams of Fig. 1(a,b), we numerically integrated Eqs. (1) in periodic square domains of size $L=512$. Regardless of the sign of $\bar{\chi}$, the system is disordered at high noise and transitions to a symmetry broken phase with nonzero global polar orders $\bm{\Pi}^{\textsl{{s}}}=\langle\hat{\bm{u}}(\theta^{\textsl{{s}}}_{k})\rangle_{k}$ when $\eta$ decreases. For all values of $\Delta\chi$, this transition is mediated by a phase-coexistence region, reminiscent of the microphase separation scenario in the single-species VM [46, 47].
@@
Mutually aligning populations.— For $\bar{\chi}>0$ at the onset of order, both species self-organize into dense bands propagating through a dilute, gaseous background. As shown in Fig. 1(c) and in the Supplementary Material [48], bands are mostly composed of a single species and are systematically arranged in pairs, with particles with the lower interspecies coupling (here b, in blue) at the rear. Such pattern synchronization by nonreciprocal alignment—also at play in discrete symmetry flocks [34]—can be understood by considering a pair of a and b bands traveling with the same velocity $v_{\rm b}^{\textsl{{a}}}=v_{\rm b}^{\textsl{{b}}}$, as shown in Fig. 1(c). Increasing the inter-band separation, particles interact with fewer neighbors of the other species, causing the local polarization in the a band to decay faster than that of the b band (since $\chi^{\textsl{{a}}\textsl{{b}}}>\chi^{\textsl{{b}}\textsl{{a}}}$). As the velocity of a band is proportional to the local polar order (from Eq. (1a)), this leads to $v_{\rm b}^{\textsl{{a}}}<v_{\rm b}^{\textsl{{b}}}$, causing the b band to catch up. With a similar reasoning, we find that reducing the inter-band separation yields $v_{\rm b}^{\textsl{{a}}}>v_{\rm b}^{\textsl{{b}}}$, such that the initial configuration is stabilized by nonreciprocity. The reversed configuration with the b band in front is, in turn, always unstable.
@@
To determine the characteristic densities $\rho^{\textsl{{a}}}_{*}$ and $\rho^{\textsl{{b}}}_{*}$ within the band, we use the position of the right and left peak of the a and b density histograms, respectively, as illustrated in Fig. 2(c). The band width $W$ is then defined from the half maximum of time-averaged a density profiles. Performing simulations in systems of increasing size, we show in Fig. 2(d) that $W$ grows linearly with $L$. Moreover, $\rho^{\textsl{{a}}}_{*}$ exhibits a finite-size behaviour which we best fit as an algebraic convergence to an asymptotic value: $\rho^{\textsl{{a}}}_{*}\simeq\rho^{\textsl{{a}}}_{\infty}-AL^{-\nu}$ with $\nu\simeq\tfrac{1}{2}$ [Fig. 2(e)]. These measures indicate that, up to strong finite size effects, the single-band phase displays characteristics of a phase-separated state in which both species spontaneously demix, despite the absence of repulsive interactions. Figures 2(f,g) illustrate how the properties of this phase vary with the amount of nonreciprocity. Namely, for $\Delta\chi$ near the homogeneous flocking phase, the band occupies most of the system and induces weak density modulations. Increasing $\Delta\chi$ rapidly enhances demixing [note the vertical logarithmic scale in Fig. 2(f)], such that for $\Delta\chi\lesssim\bar{\chi}$, the densities $\rho^{\textsl{{a}},\textsl{{b}}}_{*}$ are orders of magnitude apart and $W$ takes values close to zero. Remarkably, the resulting thin and dense band exists even for values of $\chi^{\textsl{{b}}\textsl{{a}}}=\bar{\chi}-\Delta\chi$ as low as $10^{-2}$, while it naturally disappears when $\Delta\chi=\bar{\chi}$.
@@
Figure 2: Characterisation of the single band phase. (a,b) $\rho^{\textsl{{a}}}$(a) and $\rho^{\textsl{{b}}}$(b) profiles averaged along the band and over time for different system sizes. Inset of (b): the component of the local polarity profiles along the global order. (c) Density histograms at $L=1024$. The band densities $\rho^{\textsl{{a}},\textsl{{b}}}_{*}$ (vertical dashed lines) are evaluated from the position of the relevant peaks. (d,e) Band width (d) and density (e) as functions of system size. Symbols show the numerical data and solid lines indicate fits (see text). (f,g) The densities in the band (f) and its normalized width (g) as functions of $\Delta\chi$ for various system sizes. The horizontal grey line in (f) indicates $\rho_{0}$. In all panels, $\eta=0.3$ and for (a-e) $\Delta\chi=0.4$.
@@
Mutually anti-aligning populations.— For $\bar{\chi}<0$ and low nonreciprocity, the coexistence phase at the onset of order may consist of trains of bands moving in opposite directions [Fig. 1(e)]. These apolar bands appear in equal number and are regularly spaced for $\Delta\chi=0$ [48], while their structure becomes irregular at higher nonreciprocity. Consistently with results reported for $\Delta\chi=0$ [41], these configurations coexist in a large part of the coexistence phase with polar arrangements of bands of alternating species that travel in the same direction [Fig. 1(f)]. We do not observe dynamical transitions between polar and apolar patterns, which are selected by the initial condition. Apolar bands occur more frequently at low nonreciprocity and noise, while increasing $\Delta\chi$, polar bands arise more often [48]. In fact, the latter become the only observed configuration for sufficiently large $\eta$ and $\Delta\chi$ [purple region in Fig. 1(b)].
@@
Similarly to aligning populations, the homogeneous liquid with anti-aligned mean polar orders $\langle\bm{\Pi}^{\textsl{{a}},\textsl{{b}}}\rangle_{t}$ found at $\Delta\chi=0$ is only stable up to a finite amount of nonreciprocity. Beyond a $\Delta\chi$ threshold, the two species self-organize in a laning pattern consisting of alternating domains with opposite polarities [Fig. 1(h) and Supplementary movie (SMov) 1]. This pattern is highly dynamic, as the lanes persistently rotate and drift over long timescales (SMov 2), while they also intermittently break and reform. Increasing $\Delta\chi$, breaking events become more frequent, eventually preventing the pattern from repairing itself. At sufficiently high $\Delta\chi$, the populations demix by arranging into polar clusters mainly made of a single species that move in a chaotic manner and lack macroscopic spatial order [Fig. 1(g) and SMov 3]. The lane breaking events, which are associated with a sharp decay of polar order for both species [see Fig. A1(a) in Appendix A], also happen more often in larger systems (SMov 4). As a result, the system becomes progressively disordered as revealed by the distributions of $|\bm{\Pi}^{\textsl{{s}}}|$ shown in Fig. A1(b). We therefore expect the laning pattern to eventually disappear in the thermodynamic limit, resulting in a disordered polar cluster phase.
@@
Hydrodynamic theory—To gain a theoretical understanding of the above findings, we coarse-grained the microscopic model (1) into a continuous theory. Because of its success at qualitatively capturing the physics of the single-species VM, we followed the Boltzmann-Ginzburg-Landau approach [49, 50, 51]. As detailed in Appendix B, we first write the Boltzmann equation governing the evolution of the single-body distribution function $f^{\textsl{{s}}}(\bm{r},\theta,t)$ associated with species s. By expanding $f^{\textsl{{s}}}$ in angular Fourier modes, we then obtain infinite coupled hierarchies of nonlinear equations. The zeroth and first modes correspond to the local particle densities ($\rho^{\textsl{{a}},\textsl{{b}}}$) and polarities ($\bm{p}^{\textsl{{a}},\textsl{{b}}}$), respectively, which are the physical fields of interest. Using a standard scaling ansatz describing the relative amplitude of the modes and their spatio-temporal variations at the onset of order, we truncate and close the hierarchies at the first nontrivial order. The resulting hydrodynamic equations for $\rho^{\textsl{{s}}}$ and $\bm{p}^{\textsl{{s}}}$ read
@@
where $F^{\textsl{{s}}}_{\textsl{{u}}}\equiv\alpha^{\textsl{{s}}}_{\textsl{{u}}}(\rho_{0}^{\textsl{{a}}},\rho_{0}^{\textsl{{b}}})-\sum_{\textsl{{v}},\textsl{{w}}}\xi^{\textsl{{s}}}_{\textsl{{u}}\textsl{{v}}\textsl{{w}}}s_{0}^{\textsl{{v}}}s_{0}^{\textsl{{w}}}\cos(\Delta\theta_{\textsl{{v}}\textsl{{w}}})$ and $\Delta\theta_{\textsl{{s}}\textsl{{u}}}\equiv\theta^{\textsl{{s}}}-\theta^{\textsl{{u}}}$. When $|\Delta\chi|<|\bar{\chi}|$, Eqs. (3) admit three types of solutions. For sufficiently high noise, $s_{0}^{\textsl{{a}},\textsl{{b}}}=0$ indicating a disordered state. Decreasing $\eta$ below a threshold value marked by the cyan lines in Figs. 3, the solutions bifurcate to an ordered state with $s^{\textsl{{a}},\textsl{{b}}}_{0}>0$, corresponding to a flocking ($\Delta\theta_{\textsl{{a}}\textsl{{b}}}=0$, $\bar{\chi}>0$) or anti-flocking ($\Delta\theta_{\textsl{{a}}\textsl{{b}}}=\pi$, $\bar{\chi}<0$) configuration  [^2] In fact, both solution generally exist for small $\bar{\chi}$ (more details in [48]). However, the flocking and anti-flocking solutions are only stable to homogeneous perturbations for $\bar{\chi}>0$ and $\bar{\chi}<0$, respectively. For the linear stability analysis with spatially-dependent perturbations, we considered only these solutions.. Since Eqs. (3) are nonlinear, we determine $s^{\textsl{{a}},\textsl{{b}}}_{0}$ numerically.
@@
To analyse the stability of these solutions, we linearise Eqs. (2) for small perturbations, which we express in Fourier space: $(\delta\rho^{\textsl{{s}}},\delta\theta^{\textsl{{s}}},\delta s^{\textsl{{s}}})(\bm{r},t)\to(\delta\hat{\rho}^{\textsl{{s}}},\delta\hat{\theta}^{\textsl{{s}}},\delta\hat{s}^{\textsl{{s}}})(\bm{q},t)$. We obtain a six-dimensional linear system $\partial_{t}\bm{\hat{}}{\bm{X}}=\mathbb{M}(\bm{q})\bm{\hat{}}{\bm{X}}$, where $\hat{\bm{X}}\equiv\begin{pmatrix}\delta\hat{\rho}^{\textsl{{a}}}&\delta\hat{\rho}^{\textsl{{b}}}&\delta\hat{\theta}^{\textsl{{a}}}&\delta\hat{\theta}^{\textsl{{b}}}&\delta\hat{s}^{\textsl{{a}}}&\delta\hat{s}^{\textsl{{b}}}\end{pmatrix}^{T}$ and the matrix $\mathbb{M}(\bm{q})$ is given in [48] for completeness. We construct the linear stability diagrams for both $\bar{\chi}>0$ [Fig. 3(a)] and $\bar{\chi}<0$ [Fig. 3(b)] by numerically evaluating the eigenvalues of $\mathbb{M}(\bm{q})$ for various values of $\eta$ and $\Delta\chi$, and wave vectors $\bm{q}$. In qualitative agreement with the VM phenomenology [21] and our microscopic model simulations, both the flocking and anti-flocking solutions are unstable at the onset of order. Similarly to single species flocks, the resulting long-wavelength instability is characterized by a growth rate ${\rm Re}(\Lambda){\simeq}|\bm{q}|^{2}$ for $|\bm{q}|\to 0$ [49, 53]. We uncover an additional long-wavelength instability, located deeper in the ordered phase, and which typically requires sufficiently high nonreciprocity. For both the flocking and anti-flocking cases this instability exhibits a distinct scaling: ${\rm Re}(\Lambda){\simeq}|\bm{q}|$ as $|\bm{q}|\to 0$ [inset of Fig. 3(b)], indicating that it originates from a new mechanism  [^3] For $\bar{\chi}>0$, Fig. 3(a) displays a third unstable region at low noises. As Eqs. (2) are derived perturbatively at the onset of order (see Appendix B), this instability arises far from the formal range of validity of the continuum model and, like for single species flocks, we believe that it is an artifact resulting from approximations made during the coarse-graining [51]..
@@
Figure 3: Phase diagrams in the $(\Delta\chi,\eta)$ plane obtained from the linear stability of the stationary homogeneous solutions of Eqs. (2) in the flocking (a) and anti-flocking (b) regimes. Stable disordered, flocking and anti-flocking regions appear in pink, blue and gray, respectively. The other regions indicate an instability, and are colored according to the orientation of the most unstable wave vector with respect to the direction of order. The solid cyan lines mark the onset of order, while the pink lines enclose regions where the matrix $\mathbb{M}_{\rm h}(\bm{q})$ in Eq. (4) predicts an instability. Inset of (b): Growth rate of the flocking (blue) and anti-flocking (red) instabilities as functions of the wave number $q=|\bm{q}|$ along the most unstable direction.
@@
To elucidate its origin, we consider the limit $|\bm{q}|\to 0$ where non-hydrodynamic modes can be enslaved. The hydrodynamic modes —for which $\partial_{t}\hat{(\cdot)}=\mathcal{O}(|\bm{q}|)$— correspond to the left eigenvectors of $\mathbb{M}(\bm{0})$ with zero eigenvalue. In the (anti-)flocking region, we find three of them: the two density fields $\delta\hat{\rho}^{\textsl{{a}},\textsl{{b}}}$, and $\bar{\theta}=s^{\textsl{{a}}}_{0}F^{\textsl{{b}}}_{\textsl{{a}}}\delta\hat{\theta}^{\textsl{{a}}}+{s^{\textsl{{b}}}_{0}}{F^{\textsl{{a}}}_{\textsl{{b}}}}\delta\hat{\theta}^{\textsl{{b}}}$. Enslaving the fast modes and keeping only leading order terms in $|\bm{q}|$ (details in Appendix C), the dynamics reduces to a three-dimensional linear system $\partial_{t}\hat{\bm{X}}_{\rm h}=\mathbb{M}_{\rm h}(\bm{q})\hat{\bm{X}}_{\rm h}$, where $\hat{\bm{X}}_{\rm h}=\begin{pmatrix}\delta\hat{\rho}^{\textsl{{a}}}&\delta\hat{\rho}^{\textsl{{b}}}&\bar{\theta}\end{pmatrix}^{T}$, and

### v2

@@ Abstract
We show that even weak nonreciprocal alignment leads to large-scale structure formation in flocking mixtures. By combining numerical simulations of a binary Vicsek model and the analysis of coarse-grained continuum equations, we demonstrate that nonreciprocity destabilizes the ordered phase formed by mutually aligning or anti-aligning species in a large part of the phase diagram. For aligning populations, this instability results in one species condensing in a single band that travels within a homogeneous liquid of the other species. When interactions are anti-aligning, both species self-assemble into polar clusters with large-scale chaotic dynamics. In both cases, the emergence of structures is accompanied by the demixing of the two species, despite the absence of repulsive interactions. Our theoretical analysis allows us to elucidate the origin of the instability, and show that it is generic to nonreciprocal flocks.
@@
When $\chi^{\textsl{{s}}\textsl{{u}}}=1$ for all $\textsl{{s}},\textsl{{u}}\in\{\textsl{{a}},\textsl{{b}}\}$, the model reduces to the classical VM, while other limiting cases such as fully anti-aligning mixtures [34], reciprocally anti-aligning populations ($\chi^{\textsl{{a}}\textsl{{a}},\textsl{{b}}\textsl{{b}}}>0$, $\bar{\chi}<0$, $\Delta\chi=0$[41, 38, 14] and self-anti-aligning species ($\chi^{\textsl{{a}}\textsl{{a}},\textsl{{b}}\textsl{{b}}}<0$) with aligning interspecies interactions ($\bar{\chi}>0$[36, 46] have been addressed in the literature. Here, we let particles of the same species align ($\chi^{\textsl{{a}}\textsl{{a}},\textsl{{b}}\textsl{{b}}}=1$) while varying the interspecies couplings $\chi^{\textsl{{a}}\textsl{{b}}}$ and $\chi^{\textsl{{b}}\textsl{{a}}}$. In the regime of strong nonreciprocity, where $|\Delta\chi|>|\bar{\chi}|$, we observe oscillatory behavior consistently with results reported in the literature [25, 33, 40]. Hereafter, however, we focus on the weakly nonreciprocal regime, where $|\Delta\chi|\leq|\bar{\chi}|$.
As we focus on the effects induced by nonreciprocity alone, we discard the possible influence of population heterogeneity [23] by setting $\eta^{\textsl{{a}},\textsl{{b}}}=\eta$ and $v_{0}^{\textsl{{a}},\textsl{{b}}}=v_{0}$  [^1] We show in [1] that our conclusions remain unaffected when introducing population heterogeneity.. We also consider equal densities $\rho^{\textsl{{a}},\textsl{{b}}}=\rho_{0}=N/L^{2}$, where $N$ denotes the size of each population and $L$ is the linear size of the square domain. Unless specified otherwise, we use $\rho_{0}=\tfrac{1}{2}$, $v_{0}=1$, $\bar{\chi}=\pm\tfrac{1}{2}$, and vary the angular noise $\eta$ and the nonreciprocity parameter $\Delta\chi$. Noting that the problem is symmetric under the exchange $\textsl{{a}}\leftrightarrow\textsl{{b}}$, we further restrict the analysis to $\Delta\chi\geq 0$. To construct the phase diagrams of Fig. 1(a,b), we numerically integrated Eqs. (1) in periodic square domains of size $L=512$. Regardless of the sign of $\bar{\chi}$, the system is disordered at high noise and transitions to a symmetry broken phase with nonzero global polar orders $\bm{\Pi}^{\textsl{{s}}}=\langle\hat{\bm{u}}(\theta^{\textsl{{s}}}_{k})\rangle_{k}$ when $\eta$ decreases. For all values of $\Delta\chi$, this transition is mediated by a phase-coexistence region, reminiscent of the microphase separation scenario at play in the single-species VM [27, 51].
@@
Mutually aligning populations.— For $\bar{\chi}>0$ at the onset of order, both species self-organize into dense bands propagating through a dilute, gaseous background. As shown in Fig. 1(c) and in the Supplementary Material [1], bands are mostly composed of a single species and are systematically arranged in pairs, with particles having the lower interspecies coupling (here b, in blue) at the rear. Such spatial synchronization by nonreciprocal alignment—also reported for discrete symmetry flocks [40]—can be qualitatively understood from a simple argument detailed in Appendix A.
@@
To determine the characteristic densities $\rho^{\textsl{{a}}}_{*}$ and $\rho^{\textsl{{b}}}_{*}$ within the band, we use the position of the right and left peak of the a and b density histograms, respectively, as illustrated in Fig. 2(c). The band width $W$ is then defined from the half maximum of time-averaged a density profiles. Performing simulations in systems of increasing size, we show in Fig. 2(d) that $W$ grows linearly with $L$. Moreover, $\rho^{\textsl{{a}}}_{*}$ exhibits an unusually slow convergence to its asymptotic value, which we best fit as a power law: $\rho^{\textsl{{a}}}_{*}\simeq\rho^{\textsl{{a}}}_{\infty}-AL^{-\nu}$ with $\nu\simeq\tfrac{1}{2}$ [Fig. 2(e)]. These measures indicate that, up to strong finite size effects, the single-band phase displays characteristics of a phase-separated state in which both species spontaneously demix, despite the absence of repulsive interactions. Figures 2(f,g) illustrate how the properties of this phase vary with the amount of nonreciprocity. For small $\Delta\chi$, the width of the band $W$ is comparable to the system size $L$, while it induces weak density modulations. As $L$ grows, the homogeneous flocking phase becomes more susceptible to the presence of nonreciprocity, such that the demixed phase is found for increasingly smaller values of $\Delta\chi$, down to $\Delta\chi=0.07$ for $L=8192$ ($N=67\!\cdot\!10^{6}$ particles). Despite our computational efforts, we were unable to determine whether the transition from a homogeneous flock to the demixed phase arises at a finite value of $\Delta\chi$, or if the latter can survive an infinitesimal amount of nonreciprocity. Increasing $\Delta\chi$ rapidly enhances demixing [note the vertical logarithmic scale in Fig. 2(f)], such that for $\Delta\chi\lesssim\bar{\chi}$, the densities $\rho^{\textsl{{a}},\textsl{{b}}}_{*}$ are orders of magnitude apart and $W$ takes values close to zero. Remarkably, the resulting thin and dense band exists even for values of $\chi^{\textsl{{b}}\textsl{{a}}}=\bar{\chi}-\Delta\chi$ as low as $10^{-2}$, while it naturally disappears when $\Delta\chi=\bar{\chi}$.
@@
Figure 2: Characterisation of the single band phase. (a,b) $\rho^{\textsl{{a}}}$(a) and $\rho^{\textsl{{b}}}$(b) profiles averaged along the band and over time for different system sizes. Inset of (b): the component of the local polarity profiles along the global order. (c) Density histograms at $L=1024$. $\rho^{\textsl{{a}}}_{*}$ is shown with a vertical dashed line, while $\rho^{\textsl{{b}}}_{*}$ is defined as the position of left peak of the blue histogram. (d,e) Band width (d) and density (e) as functions of system size. Symbols show the numerical data and solid lines indicate fits (see text). (f,g) The densities in the band (f) and its normalized width (g) as functions of $\Delta\chi$ for various system sizes. The horizontal grey line in (f) indicates $\rho_{0}$. In all panels, $\eta=0.3$ and for (a-e) $\Delta\chi=0.4$.
@@
Mutually anti-aligning populations.— For $\bar{\chi}<0$ and low nonreciprocity, the coexistence phase at the onset of order may consist of trains of bands moving in opposite directions [Fig. 1(e)]. In these apolar arrangements, bands appear in equal number and are regularly spaced for $\Delta\chi=0$ [1], while their structure becomes irregular at higher nonreciprocity. Consistently with results reported for $\Delta\chi=0$ [14], these configurations coexist in a large part of the coexistence phase with polar arrangements of bands of alternating species that travel in the same direction [Fig. 1(f)]. We do not observe dynamical transitions between polar and apolar patterns, which are selected by the initial condition. Apolar configurations occur more frequently at low nonreciprocity and noise, while increasing $\Delta\chi$, polar trains arise more often [1]. In fact, the latter become the only observed configuration for sufficiently large $\eta$ and $\Delta\chi$ [purple region in Fig. 1(b)].
@@
Similarly to aligning populations, the homogeneous liquid with anti-aligned mean polar orders $\langle\bm{\Pi}^{\textsl{{a}},\textsl{{b}}}\rangle_{t}$ found at $\Delta\chi=0$ is only stable up to a finite amount of nonreciprocity. Beyond a $\Delta\chi$ threshold, the two species self-organize in a laning pattern consisting of alternating domains with opposite polarities [Fig. 1(h) and Supplementary movie SMov 1]. This pattern is highly dynamic, as the lanes persistently rotate and drift over long timescales (SMov 2), while they also intermittently break and reform. Increasing $\Delta\chi$, breaking events become more frequent, eventually preventing the pattern from repairing itself. At sufficiently high $\Delta\chi$, the populations demix by arranging into polar clusters mainly made of a single species, which move in a chaotic manner and lack macroscopic spatial order [Fig. 1(g) and SMov 3]. The lane breaking events, which are associated with a sharp decay of polar order for both species [see Fig. B1(a) in Appendix B], happen more often in larger systems (SMov 4). As a result, the system becomes progressively disordered as revealed by the distributions of $|\bm{\Pi}^{\textsl{{s}}}|$ shown in Fig. B1(b). We therefore expect the laning pattern to eventually disappear in the thermodynamic limit, resulting in a disordered polar cluster phase.
@@
Hydrodynamic theory—To gain a theoretical understanding of the above findings, we coarse-grained the microscopic model (1) into a continuous theory. Because of its success at qualitatively capturing the physics of the single-species VM, we followed the Boltzmann-Ginzburg-Landau approach [6, 7, 47]. As detailed in Appendix C, we first write the Boltzmann equation governing the evolution of the single-body distribution function $f^{\textsl{{s}}}(\bm{r},\theta,t)$ associated with species s. By expanding $f^{\textsl{{s}}}$ in angular Fourier modes, we then obtain infinite coupled hierarchies of nonlinear equations. The zeroth and first modes correspond to the local particle densities ($\rho^{\textsl{{a}},\textsl{{b}}}$) and polarities ($\bm{p}^{\textsl{{a}},\textsl{{b}}}$), respectively, which are the physical fields of interest. Using a standard scaling ansatz describing the relative amplitude of the modes and their spatio-temporal variations at the onset of order, we truncate and close the hierarchies at the first nontrivial order. The resulting hydrodynamic equations for $\rho^{\textsl{{s}}}$ and $\bm{p}^{\textsl{{s}}}$ read
@@
where $F^{\textsl{{s}}}_{\textsl{{u}}}\equiv\alpha^{\textsl{{s}}}_{\textsl{{u}}}(\rho_{0}^{\textsl{{a}}},\rho_{0}^{\textsl{{b}}})-\sum_{\textsl{{v}},\textsl{{w}}}\xi^{\textsl{{s}}}_{\textsl{{u}}\textsl{{v}}\textsl{{w}}}s_{0}^{\textsl{{v}}}s_{0}^{\textsl{{w}}}\cos(\Delta\theta_{\textsl{{v}}\textsl{{w}}})$ and $\Delta\theta_{\textsl{{s}}\textsl{{u}}}\equiv\theta^{\textsl{{s}}}-\theta^{\textsl{{u}}}$. When $|\Delta\chi|<|\bar{\chi}|$, Eqs. (3) admit three types of solutions. For sufficiently high noise, $s_{0}^{\textsl{{a}},\textsl{{b}}}=0$ indicating a disordered state. Decreasing $\eta$ below a threshold value marked by the cyan lines in Figs. 3, the solutions bifurcate to an ordered state with $s^{\textsl{{a}},\textsl{{b}}}_{0}>0$, corresponding to a flocking ($\Delta\theta_{\textsl{{a}}\textsl{{b}}}=0$, $\bar{\chi}>0$) or anti-flocking ($\Delta\theta_{\textsl{{a}}\textsl{{b}}}=\pi$, $\bar{\chi}<0$) configuration.
@@
To analyse the stability of these solutions, we linearise Eqs. (2) for small perturbations, which we express in Fourier space: $(\delta\rho^{\textsl{{s}}},\delta\theta^{\textsl{{s}}},\delta s^{\textsl{{s}}})(\bm{r},t)\to(\delta\hat{\rho}^{\textsl{{s}}},\delta\hat{\theta}^{\textsl{{s}}},\delta\hat{s}^{\textsl{{s}}})(\bm{q},t)$. We obtain a six-dimensional linear system $\partial_{t}\bm{\hat{}}{\bm{X}}=\mathbb{M}(\bm{q})\bm{\hat{}}{\bm{X}}$, where $\hat{\bm{X}}\equiv\begin{pmatrix}\delta\hat{\rho}^{\textsl{{a}}}&\delta\hat{\rho}^{\textsl{{b}}}&\delta\hat{\theta}^{\textsl{{a}}}&\delta\hat{\theta}^{\textsl{{b}}}&\delta\hat{s}^{\textsl{{a}}}&\delta\hat{s}^{\textsl{{b}}}\end{pmatrix}^{T}$ and the matrix $\mathbb{M}(\bm{q})$ is given in [1] for completeness. We construct the linear stability diagrams for both $\bar{\chi}>0$ [Fig. 3(a)] and $\bar{\chi}<0$ [Fig. 3(b)] by numerically evaluating the eigenvalues of $\mathbb{M}(\bm{q})$ for various values of $\eta$ and $\Delta\chi$, and wave vectors $\bm{q}$. In qualitative agreement with the VM phenomenology [13] and our microscopic model simulations, both the flocking and anti-flocking solutions are unstable at the onset of order. Similarly to single species flocks, the resulting long-wavelength instability is characterized by a growth rate ${\rm Re}(\Lambda){\simeq}|\bm{q}|^{2}$ for $|\bm{q}|\to 0$ [6, 12]. We uncover an additional long-wavelength instability, located deeper in the ordered phase, and which typically requires sufficiently high nonreciprocity. For both the flocking and anti-flocking cases this instability exhibits a distinct scaling: ${\rm Re}(\Lambda){\simeq}|\bm{q}|$ as $|\bm{q}|\to 0$ [inset of Fig. 3(b)], indicating that it originates from a new mechanism  [^2] For $\bar{\chi}>0$, Fig. 3(a) displays a third unstable region at low noises. As Eqs. (2) are derived perturbatively at the onset of order, this instability arises far from the formal range of validity of the continuum model and, like for single species flocks, we believe that it is an artifact resulting from approximations made during the coarse-graining [1, 47]..
@@
Figure 3: Phase diagrams in the $(\Delta\chi,\eta)$ plane obtained from the linear stability of the stationary homogeneous solutions of Eqs. (2) in the flocking (a) and anti-flocking (b) regimes. Stable disordered, flocking and anti-flocking regions appear in green, blue and gray, respectively. The other regions indicate an instability, and are colored according to the orientation of the most unstable wave vector with respect to the direction of order. The solid cyan lines mark the onset of order, while the pink lines enclose regions where the matrix $\mathbb{M}_{\rm h}(\bm{q})$ in Eq. (4) predicts an instability. Inset of (b): Growth rate of the flocking (blue) and anti-flocking (red) instabilities as functions of the wave number $q=|\bm{q}|$ along the most unstable direction.
@@
To elucidate its origin, we consider the limit $|\bm{q}|\to 0$ where non-hydrodynamic modes can be enslaved. The hydrodynamic modes —for which $\partial_{t}\hat{(\cdot)}=\mathcal{O}(|\bm{q}|)$— correspond to the left eigenvectors of $\mathbb{M}(\bm{0})$ with zero eigenvalue. In the (anti-)flocking region, we find three of them: the two density fields $\delta\hat{\rho}^{\textsl{{a}},\textsl{{b}}}$, and $\bar{\theta}=s^{\textsl{{a}}}_{0}F^{\textsl{{b}}}_{\textsl{{a}}}\delta\hat{\theta}^{\textsl{{a}}}+{s^{\textsl{{b}}}_{0}}{F^{\textsl{{a}}}_{\textsl{{b}}}}\delta\hat{\theta}^{\textsl{{b}}}$. Enslaving the fast modes and keeping only leading order terms in $|\bm{q}|$ (details in Appendix D), the dynamics reduces to a three-dimensional linear system $\partial_{t}\hat{\bm{X}}_{\rm h}=\mathbb{M}_{\rm h}(\bm{q})\hat{\bm{X}}_{\rm h}$, where $\hat{\bm{X}}_{\rm h}=\begin{pmatrix}\delta\hat{\rho}^{\textsl{{a}}}&\delta\hat{\rho}^{\textsl{{b}}}&\bar{\theta}\end{pmatrix}^{T}$, and